# Kernel-based optimization of measurement operators for quantum reservoir computers

### v1

@@ IV.1 Image classification task
We summarize this section with some common findings: (i) Unless one uses a complete operator basis or techniques like temporal multiplexing [31], the primal optimization is generally inferior to the kernel-based approach: in the latter, it suffices to restrict the readout to Pauli-$Z$ observables enhanced by 2nd-degree monomials to reach the classical SVM accuracy $\mathcal{A}_{\text{CSVM}}$. The fundamental reason is of course that the kernel-based method implicitly renders an optimized unitary to be used as pre-rotation before measuring. (ii) Performing the kernel-based optimization only over a single class can significantly reduce the computational load without sacrificing accuracy, provided the QELM is retrained for all classes. (iii) Low order monomial in the readout (especially squares) are generally helpful to close the quantum-to-classical accuracy gap (depending on the specific encoding).
@@ IV.2 Time series prediction
We turn to autoregressive prediction tasks, where, given a time series $\{\mathbf{z}_{t}\}_{t=0}^{T}$, the QRC is trained to predict the next step $\hat{\mathbf{z}}_{t+1}=f(\mathbf{z}_{t})$ [31, 33, 52]. A complete forecasted trajectory is generated by repeatedly applying the prediction function, i.e., $\hat{\mathbf{z}}_{t+1}=f(\hat{\mathbf{z}}_{t})$. As was the case for static data, enhancing the readout with monomials typically improves forecasting capability [48, 37]. However, they obscure a clear assessment of the performance of the quantum parts of the model, and we therefore consider here only setups without hand-engineered features in the readout. As a performance measure, we consider, beside the root-mean-square training error $E_{\text{train}}=\left[\frac{1}{P}\sum_{t=0}^{P-1}\|\mathbf{z}_{t+1}-f(\mathbf{z}_{t})\|^{2}\right]^{1/2}$, the forecast horizon
@@ IV.2.2 QRCs with internal memory
Using memory states significantly enhances prediction accuracy both for the primal and dual optimization method (Fig. 3), yielding forecast horizons of $T_{\text{fch}}/T_{L}\approx 10$ (which is comparable to recent literature results [37]). Notably, using the optimal operator $M^{*}$ directly leads in this case to a significantly lower forecast horizon and higher training error ($E_{\text{train}}\approx 2\times 10^{-6}$) compared to the Pauli decomposition or diagonalization methods ($E_{\text{train}}\approx 4\times 10^{-8}$). We attribute this to the fact that we re-fit the operator subsets to the data, which enables the model to attain a higher expressivity compared to using just a single $M^{*}$ per output. However, this appears to be specific to current QRC setup, as we do not observe such strong discrepancy in the other benchmarks (cf. Figs. 2 and 4).
@@
$z(t)=\sum_{k=1}^{n}\left[a_{k}\cos(k\omega_{0}t)+b_{k}\sin(k\omega_{0}t)\right]$ (37)
Learning this time series with a classical linear autoregressive model requires a memory capacity of $2n$ delays. Due to the nonlinearity generated by the encoding and the ancilla coupling of a QRC like Eq. 18, this bound is only an upper limit in terms of actual number of required linear encoding features. Specifically, the number of ancilla qubits $N_{A}$ required to store the delays is much less than $2n$. While a detailed study will be presented elsewhere, we find here that $N_{A}=3$ and a well tuned memory capacity (via the TFIM parameters) suffices to learn $n\sim 10$ frequencies. We use amplitude encoding [Eq. 29] and a resolution $\Delta t=0.2$  [^3] The valid region of $\Delta t$ follows from the requirements that (i) the fading memory window covers a full period and (ii) the sampling rate larger than $2f_{\text{max}}$ (Nyquist criterion). to feed the time samples $z_{n}=z(n\Delta t)$ into the QRC.
@@
with chaotic-regime parameters $\theta=17$, $\beta=0.2$, $\gamma=0.1$, $n=10$. We evolve Eq. 38 numerically with a time step $\Delta t=1$. Samples $u_{n}=u(n\Delta t)$ are scaled to the interval $[0,0.3]$ and injected with the amplitude encoding variant of Eq. 28 into a QRC with TFIM-based unitary and $N_{A}=3$ ancilla qubits. As seen in Fig. 4(b), the optimized QRC can forecast the Mackey-Glass time series up to $\geq 250$ time steps. The Pauli-decomposition with slightly reduced observable count reaches even to $\sim 300$ time steps, possibly due to numerical stabilization effects not present when using the full operator $M^{*}$ or the Pauli-$Z$ decomposition. Note that the forecast horizon is necessarily finite due to the intrinsic Lyapunov instability of the Mackey-Glass system, with Lyapunov exponents reported of $\lambda_{1}\simeq 6\times 10^{-3}$ [60].
@@ V Summary
1. Data Input pairs $\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{P}$ Time series $\{\mathbf{z}_{t}\}_{t=0}^{T-1}$; Targets $y_{t}=\mathbf{z}_{t+1}$
@@ 2. Encoding Map each $\mathbf{x}_{i}$ to state: To obtain $\rho(x_{t})$, evolve for $\tau=t-L,\ldots,t$:
3. Kernel Gram matrix: $K_{ij}=K(\mathbf{x}_{i},\mathbf{x}_{j})=\mathrm{tr}(\rho(\mathbf{x}_{i})\rho(\mathbf{x}_{j}))$
@@
5. Optimal Observable $M^{*}=\sum_{k}^{P}\alpha_{k}^{*}\rho_{k}$
6. Prediction For new input $\mathbf{x}$: For new input $\mathbf{z}_{t}$:
$f(\mathbf{x})=\mathrm{tr}(M^{*}\rho(\mathbf{x}))$ $f(\mathbf{z}_{t})=\mathrm{tr}(M^{*}\rho_{t}(\mathbf{z}_{t}))=\sum_{t^{\prime}}\alpha_{t^{\prime}}^{*}K(x_{t^{\prime}},\mathbf{z}_{t})$
$\sum_{i}\alpha_{i}^{*}K(\mathbf{x}_{i},\mathbf{x})$
@@ 6. Prediction For new input $\mathbf{x}$: For new input $\mathbf{z}_{t}$:
Table 3: Algorithm flow for the kernel-based (dual) optimization of the measurement operator in stateless (memoryless) and stateful (recurrent) QRCs.
@@
We have provided a framework, based on the connection between quantum kernel methods and supervised quantum ML models [19], to optimize the measurement operator in quantum reservoir computers (QRCs). Our methods is applicable to both stateless (QELM) and stateful models, as summarized by the algorithm flow chart in Table 3. Since the essence of the method does not rely on the QRC-specific form (2) of the output function, insights gained from this work can be transferred to other quantum ML models with a well-defined (and already trained) feature map $\mathbf{x}\mapsto\rho(\mathbf{x})$, enhancing previous discussions of primal-dual approaches in QML [22, 61].
@@
Exact implementation of the optimal measurement operator $M^{*}$ on an $N$-qubit system requires measuring all $2^{N}$ $Z$-strings (or, equivalently, measuring in the computational basis) and a pre-measurement rotation of the quantum state by a global (data-dependent) unitary. This unitary can be understood to capture the outcome otherwise achieved by training a parametrized QML model with fixed measurement operator. Interestingly, we find that using operator subsets (e.g., Pauli $Z$-strings) obtained from decomposing $M^{*}$ typically improves forecasting performance compared to using the single operator $M^{*}$ directly. This could be related to the emergence of numerical instabilities or enhanced expressivity in certain QRC setups and should be investigated further.
@@
Having determined an optimal measurement operator raises the question whether and to which extent the prediction performance of a QRC can be further improved. As will be analyzed in detail in a forthcoming work, in QML models with a time-independent encoding layer, the available features are fully determined by the encoding  [^4] For the product encoding schemes considered in Section III.2, the features are essentially polynomials of certain basis functions (e.g., Fourier exponentials or monomials in the input data).. Subsequently, the features are linearly processed by the quantum channels. This manifests as information scrambling and can reduce the prediction performance of QML models. The optimal measurement operator allows to reverse this effect via the implicit unitary it entails [see Eq. 14]. This is reflected by the fact that the trained QELM predictor (11) is invariant to global conjugation  [^5] The situation of a QRC with memory [Eq. 18] is more subtle due to the information loss represented by the trace-out operation. Once the optimal measurement operator has been determined, optimization of the reservoir unitary is not necessary anymore. Instead, meaningful optimizations would pertain to the encoding, memory depth and possibly hand-engineered readout features. Alternatively to determining the optimal measurement operator, one may use temporal multiplexing [31, 32] or measuring over a complete measurement basis (full state tomography).
@@
We have numerically demonstrated the feasibility of the optimization for QRCs with $N\lesssim 10$ qubits, noting that, for our considered datasets, it suffices to use a representative subset of the full training samples to evaluate the kernel matrix. Application to larger systems is instead bottlenecked by memory consumption of the $4^{N}$-dimensional operator matrix and requires further methodological developments. Another avenue for future work is a closer look into generalization capability entailed by the optimal measurement operator, since both primal and dual optimization methods provide only error bounds on the training set [64, 12].
@@ Appendix A QELM as a kernel method
The representer property holds directly in this operator space [19, 68]. Decompose any $M$ as $M=M_{\parallel}+M_{\perp}$, where $M_{\parallel}\in\mathrm{span}{\rho(\mathbf{x}_{1}),\ldots,\rho(\mathbf{x}_{P})}$ and $M_{\perp}$ is Hilbert–Schmidt orthogonal to this span. Then $\mathrm{tr}(M_{\perp}\rho(\mathbf{x}_{i}))=0$ for all $i$, so the data-fit term in Eq. 40 depends only on $M_{\parallel}$, while $(|M|^{\text{HS}})^{2}=(|M_{\parallel}|^{\text{HS}})^{2}+(|M_{\perp}|^{\text{HS}})^{2}$. Replacing $M$ by $M_{\parallel}$ strictly decreases the objective unless $M_{\perp}=0$. Therefore, there exists an optimal observable $M^{*}$ in the span of the training data, i.e.,
@@
$\min_{\boldsymbol{\alpha}\in\mathbb{R}^{P}}\sum_{i=1}^{P}\mathcal{L}\big([\mathbf{K}\boldsymbol{\alpha}]_{i},y_{i}\big)+\frac{\lambda}{2}\boldsymbol{\alpha}^{\top}\mathbf{K}\boldsymbol{\alpha}$ (44)
@@
$\min_{\boldsymbol{\alpha}\in\mathbb{R}^{P}}\frac{1}{2}\,|\mathbf{K}\boldsymbol{\alpha}-\mathbf{y}|_{2}^{2}+\frac{\lambda}{2}\boldsymbol{\alpha}^{\top}\mathbf{K}\boldsymbol{\alpha}.$ (45)
@@
$\max_{\boldsymbol{\beta}\in\mathbb{R}^{P}}\sum_{i=1}^{P}\beta_{i}-\frac{1}{2}\sum_{i,j=1}^{P}\beta_{i}\beta_{j}y_{i}y_{j}K_{ij}\quad\text{subject to}\quad 0\leq\beta_{i}\leq\frac{1}{\lambda},$ (46)
with $\alpha_{i}^{*}=y_{i}\beta_{i}^{*}$. For the latter loss, the solution coefficients $\boldsymbol{\alpha}^{*}$ are not available in closed form, but can be computed numerically.
@@
Expanding $M^{*}$ in terms of an orthonormal operator basis $\{M_{k}\}$ yields $M^{*}=\sum_{k}w_{k}M_{k}$ with
$w_{k}^{*}=\mathrm{tr}(M_{k}M^{*})=\sum_{i=1}^{P}\alpha_{i}^{*}\mathrm{tr}(M_{k}\rho(\mathbf{x}_{i}))$ (49)
Thus, taking
@@
$f^{*}(\mathbf{x})=\phi(\mathbf{x})^{T}\mathbf{w}^{*},\qquad\mathbf{w}^{*}=\Phi^{T}\boldsymbol{\alpha}^{*}$ (51)
Using the HS completeness identity $\sum_{k}\mathrm{tr}(M_{k}A)\mathrm{tr}(M_{k}B)=\mathrm{tr}(AB)$ for any operators $A,B$, it is easy to see that, indeed, $f^{*}(\mathbf{x})=\sum_{k,i}\mathrm{tr}(M_{k}\rho(x))\alpha_{i}\mathrm{tr}(M_{k}\rho(\mathbf{x}_{i}))=\sum_{i}\alpha_{i}K(\mathbf{x},\mathbf{x}_{i})$.
Consider now the case where $\{M_{k}\}_{k=1}^{m}$ is not a complete basis, i.e., there is no exact $\mathbf{w}$ with $\sum_{k}w_{k}M_{k}=M^{*}$. The natural generalization of Eq. 49 is to find $\mathbf{w}$ that minimize the Hilbert-Schmidt distance $\|\sum_{k}w_{k}M_{k}-M^{*}\|_{\text{HS}}=\mathrm{tr}\big[(\sum_{k}w_{k}M_{k}-M^{*})^{2}\big]$. Solving this optimization problem yields $\mathbf{w}=G^{+}\Phi^{T}\alpha^{*}$, or component-wise
$w_{k}=\sum_{l=1}^{m}(G^{+})_{kl}\sum_{j=1}^{P}\alpha_{j}^{*}\mathrm{tr}(M_{l}\rho(\mathbf{x}_{j})),\qquad G_{kl}\equiv\mathrm{tr}(M_{k}M_{l})$ (52)
If $\{M_{k}\}$ is HS-orthogonal, then $G=\mathbbm{1}$ and Eq. 52 reduces to Eq. 49.

### v2

@@ IV.1 Image classification task
We summarize this section with some common findings: (i) Unless one uses a complete operator basis or techniques like temporal multiplexing [32], the primal optimization is generally inferior to the kernel-based approach: in the latter, it suffices to restrict the readout to Pauli-$Z$ observables enhanced by 2nd-degree monomials to reach the classical SVM accuracy $\mathcal{A}_{\text{CSVM}}$. The fundamental reason is of course that the kernel-based method implicitly renders an optimized pre-rotation before measurement. (ii) Performing the kernel-based optimization only over a single class can significantly reduce the computational load without sacrificing accuracy, provided the QELM is retrained for all classes. (iii) Low order monomial in the readout (especially squares) are generally helpful to close the quantum-to-classical accuracy gap (depending on the specific encoding).
@@ IV.2 Time series prediction
We turn to autoregressive prediction tasks [32, 34, 53], where, given a time series $\{\mathbf{z}_{t}\}_{t=0}^{T}$, the QRC is trained to predict the next step $\hat{\mathbf{z}}_{t+1}=f(x_{t})$ (with the prescription (3) where appropriate). A complete forecasted trajectory is generated by repeatedly applying the prediction function, i.e., $\hat{\mathbf{z}}_{t+1}=f(\hat{x}_{t})$. As was the case for static data, enhancing the readout with monomials typically improves forecasting capability [49, 38]. However, they obscure a clear assessment of the performance of the quantum parts of the model, and we therefore consider here only setups without hand-engineered features in the readout. As a performance measure, we consider, beside the root-mean-square training error $E_{\text{train}}=\left[\frac{1}{P}\sum_{t=0}^{P-1}\|\mathbf{z}_{t+1}-f(\mathbf{z}_{t})\|^{2}\right]^{1/2}$, the forecast horizon
@@ IV.2.2 QRCs with internal memory
Using memory states significantly enhances prediction accuracy both for the primal and dual optimization method (Fig. 3), yielding forecast horizons of $T_{\text{fch}}/T_{L}\approx 10$ (which is comparable to recent literature results [38]). Notably, using the optimal operator $M^{*}$ directly leads in this case to a significantly lower forecast horizon and higher training error ($E_{\text{train}}\approx 2\times 10^{-6}$) compared to the Pauli decomposition or diagonalization methods ($E_{\text{train}}\approx 4\times 10^{-8}$). This result appears to be specific to the current QRC setup or data, as we do not observe strong discrepancies in the other benchmarks (cf. Figs. 2 and 4).
@@
$z(t)=\sum_{k=1}^{n}\left[a_{k}\cos(k\omega_{0}t)+b_{k}\sin(k\omega_{0}t)\right].$ (37)
Learning this time series with a classical linear autoregressive model requires a memory capacity of $2n$ delays [59, 60]. Due to the nonlinearity generated by the encoding and the ancilla coupling of a QRC like Eq. 18, this bound is only an upper limit in terms of actual number of required linear encoding features. Specifically, the number of ancilla qubits $N_{A}$ required to store the delays is much less than $2n$. While a detailed study will be presented elsewhere, we find here that $N_{A}=3$ and a well tuned memory capacity (via the TFIM parameters) suffices to learn $n\sim 10$ frequencies. We use amplitude encoding [Eq. 29] and a resolution $\Delta t=0.2$ to feed the time samples $z_{n}=z(n\Delta t)$ into the QRC. The valid region of $\Delta t$ follows from the requirements that (i) the fading memory window covers a full period and (ii) the sampling is rate larger than $2f_{\text{max}}$ (Nyquist criterion).
@@
with chaotic-regime parameters $\theta=17$, $\beta=0.2$, $\gamma=0.1$, $n=10$. We evolve Eq. 38 numerically with a time step $\Delta t=1$. Samples $z_{n}=z(n\Delta t)$ are scaled to the interval $[0,0.3]$ and injected with the amplitude encoding variant of Eq. 28 into a QRC with TFIM-based unitary and $N_{A}=3$ ancilla qubits. As seen in Fig. 4(b), the optimized QRC can forecast the Mackey-Glass time series up to $\geq 250$ time steps. The Pauli-decomposition with slightly reduced observable count reaches even to $\sim 300$ time steps, possibly due to numerical stabilization effects not present when using the full operator $M^{*}$ or the Pauli-$Z$ decomposition. Note that the forecast horizon is necessarily finite due to the intrinsic Lyapunov instability of the Mackey-Glass system, with Lyapunov exponents reported of $\lambda_{1}\simeq 6\times 10^{-3}$ [62].
@@ V Summary
1. Data Input pairs $\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{P}$ Time series $\{\mathbf{z}_{t}\}_{t=0}^{T}$; Targets $y_{t}=\mathbf{z}_{t+1}$
@@ 2. Encoding Map each $\mathbf{x}_{i}$ to state: To obtain $\rho(x_{t})$, evolve for $\tau=t-L,\ldots,t$:
3. Kernel Gram matrix: $K_{ij}=K(x_{i},x_{j})=\mathrm{tr}(\rho_{i}\rho_{j})$
@@
5. Optimal Observable $M^{*}=\sum_{k=1}^{P}\alpha_{k}^{*}\rho_{k}$
6. Prediction For new input $\hat{\mathbf{x}}$: For new input $\hat{x}_{t}=(\hat{\mathbf{z}}_{t-L},\ldots,\hat{\mathbf{z}}_{t})$:
$)f(\hat{\mathbf{x}})=\mathrm{tr}(M^{*}\rho(\hat{\mathbf{x}}$)) $f(\hat{x}_{t})=\mathrm{tr}(M^{*}\rho(\hat{x}_{t}))=\sum_{t^{\prime}}\alpha_{t^{\prime}}^{*}K(x_{t^{\prime}},\hat{x}_{t})$
[…]
@@ 5. Optimal Observable $M^{*}=\sum_{k=1}^{P}\alpha_{k}^{*}\rho_{k}$
Table 3: Algorithm flow for the kernel-based (dual) optimization of the measurement operator in stateless (memoryless) and stateful (recurrent) QRCs. Note that $\mathbf{y}\in\mathbb{R}^{P}$ is the vector of target samples and the expression for the dual coefficients $\alpha^{*}$ thus applies componentwise. In the stateful case, P=TLP=T-L is the number of time steps available for training. For details, see Appendices A and F.
@@
We have provided a framework, based on the connection between quantum kernel methods and supervised quantum ML models [20], to optimize the measurement operator in quantum reservoir computers (QRCs). This establishes a quantum kernel formulation on history space for recurrent QRCs, connecting the optimal readout observable to the kernel methods previously employed for memoryless quantum models. Our method is thus applicable to both stateless (QELM) and stateful models, as summarized by the algorithm flow chart in Table 3. Since the essence of the method does not rely on the QRC-specific form (2) of the output function, insights gained from this work can be transferred to other quantum ML models with a well-defined (and already trained) feature map $\mathbf{x}\mapsto\rho(\mathbf{x})$, enhancing previous discussions of primal-dual approaches in QML [23, 63].
@@
Exact implementation of the optimal measurement operator $M^{*}$ on an $N$-qubit system requires measuring all $2^{N}$ $Z$-strings (or, equivalently, measuring in the computational basis) and a pre-measurement rotation of the quantum state by a global (training data-dependent) unitary. This unitary can be understood to capture the outcome otherwise achieved by training a parametrized QML model with fixed measurement operator. Empirically, we find that using operator subsets (e.g., Pauli $Z$-strings) obtained from decomposing $M^{*}$ can sometimes improve forecasting performance compared to using the single operator $M^{*}$ directly. This could be related to the emergence of numerical instabilities or conditioning effects in certain QRC setups and should be investigated further.
@@
Having determined an optimal measurement operator raises the question whether and to which extent the prediction performance of a QRC can be further improved. As will be further discussed elsewhere, in QML models with a time-independent initial-state encoding strategy, the available features are fully determined by the encoding and are subsequently linearly processed by the quantum channels  [^3] For the product encoding schemes considered in Section III.2, the features are essentially polynomials of certain basis functions (e.g., Fourier exponentials or monomials in the input data).. This processing can manifest as information scrambling and reduce the prediction performance of QML models. In the case of stateless QELMs, the unitary associated with the optimal observable can undo this unitary scrambling [see Eq. 14], which is reflected by the invariance of the trained predictor (11) under global conjugation. Accordingly, for stateless QELMs with a unitary channel, optimization of the reservoir unitary can be traded against optimization of the measurement operator, up to the final measurement/readout parametrization. By contrast, for stateful QRCs with decoherence/reset, or generally for non-unitary channels, optimizing the reservoir dynamics can change the kernel and thus the attainable predictor class. Further meaningful optimizations pertain to the encoding, memory depth and possibly hand-engineered readout features. Alternatively to determining the optimal measurement operator, one may use temporal multiplexing [32, 33] or measuring over a complete measurement basis (full state tomography).
@@
We have numerically demonstrated the feasibility of the optimization for QRCs with $N\lesssim 10$ qubits, noting that, for our considered datasets, it suffices to use a representative subset of the full training samples to evaluate the kernel matrix. Application to larger systems is instead bottlenecked by memory consumption of the $4^{N}$-dimensional operator matrix and requires further methodological developments. Another avenue for future work is a closer look into generalization capability entailed by the optimal measurement operator, since both primal and dual optimization methods provide only error bounds on the training set [65, 66, 12].
@@ Appendix A QELM as a kernel method
The representer property holds directly in this operator space [20, 70]. Decompose any $M$ as $M=M_{\parallel}+M_{\perp}$, where $M_{\parallel}\in\mathrm{span}\{\rho(\mathbf{x}_{1}),\ldots,\rho(\mathbf{x}_{P})\}$ and $M_{\perp}$ is Hilbert–Schmidt orthogonal to this span. Then $\mathrm{tr}(M_{\perp}\rho(\mathbf{x}_{i}))=0$ for all $i$, so the data-fit term in Eq. 40 depends only on $M_{\parallel}$, while $(|M|^{\text{HS}})^{2}=(|M_{\parallel}|^{\text{HS}})^{2}+(|M_{\perp}|^{\text{HS}})^{2}$. Replacing $M$ by $M_{\parallel}$ strictly decreases the objective unless $M_{\perp}=0$. Therefore, there exists an optimal observable $M^{*}$ in the span of the training data, i.e.,
@@
$\min_{\boldsymbol{\alpha}\in\mathbb{R}^{P}}\sum_{i=1}^{P}\mathcal{L}\big([\mathbf{K}\boldsymbol{\alpha}]_{i},y_{i}\big)+\frac{\lambda}{2}\boldsymbol{\alpha}^{\top}\mathbf{K}\boldsymbol{\alpha}.$ (44)
@@
$\min_{\boldsymbol{\alpha}\in\mathbb{R}^{P}}\frac{1}{2}\,\|\mathbf{K}\boldsymbol{\alpha}-\mathbf{y}\|_{2}^{2}+\frac{\lambda}{2}\boldsymbol{\alpha}^{\top}\mathbf{K}\boldsymbol{\alpha},$ (45)
@@
$\max_{\boldsymbol{\beta}\in\mathbb{R}^{P}}\sum_{i=1}^{P}\beta_{i}-\frac{1}{2\lambda}\sum_{i,j=1}^{P}\beta_{i}\beta_{j}y_{i}y_{j}K_{ij}\quad\text{subject to}\quad 0\leq\beta_{i}\leq 1,$ (46)
with $\alpha_{i}^{*}=y_{i}\beta_{i}^{*}/\lambda$ (and no bias). In this case, the solution coefficients $\boldsymbol{\alpha}^{*}$ are not available in closed form, but can be computed numerically.
@@
Expanding $M^{*}$ in terms of an orthogonal operator basis $\{M_{k}\}$ with $G_{kl}\equiv\mathrm{tr}(M_{k}M_{l})=g_{k}\delta_{kl}$, yields $M^{*}=\sum_{k}w_{k}M_{k}$ with
$w_{k}^{*}=\frac{1}{g_{k}}\mathrm{tr}(M_{k}M^{*})=\frac{1}{g_{k}}\sum_{i=1}^{P}\alpha_{i}^{*}\mathrm{tr}(M_{k}\rho(\mathbf{x}_{i})).$ (49)
For the Pauli basis, $g_{k}=2^{N}$. Thus, taking
@@
$f^{*}(\mathbf{x})=\phi(\mathbf{x})^{T}\mathbf{w}^{*},\qquad\mathbf{w}^{*}=G^{-1}\Phi^{T}\boldsymbol{\alpha}^{*}.$ (51)
Using the HS completeness identity $\sum_{k,l}\mathrm{tr}(M_{k}A)(G^{-1})_{kl}\mathrm{tr}(M_{l}B)=\mathrm{tr}(AB)$ for any Hermitian operators $A,B$, it is easy to see that, indeed, $f^{*}(\mathbf{x})=\sum_{i}\alpha_{i}^{*}K(\mathbf{x},\mathbf{x}_{i})$.
Consider now the case where $\{M_{k}\}_{k=1}^{m}$ is not a complete basis, i.e., one cannot generally assume that there exists a $\mathbf{w}$ with $\sum_{k}w_{k}M_{k}=M^{*}$. The natural generalization of Eq. 49 is to find a $\mathbf{w}$ that minimizes the (squared) Hilbert–Schmidt distance $\|\sum_{k}w_{k}M_{k}-M^{*}\|_{\text{HS}}^{2}=\mathrm{tr}\big[(\sum_{k}w_{k}M_{k}-M^{*})^{2}\big]$. Solving this optimization problem yields $\mathbf{w}=G^{+}\Phi^{T}\alpha^{*}$, or component-wise
$w_{k}=\sum_{l=1}^{m}(G^{+})_{kl}\sum_{j=1}^{P}\alpha_{j}^{*}\mathrm{tr}(M_{l}\rho(\mathbf{x}_{j})).$ (52)